OpenStudy (anonymous):

I am missing something simple. How do I find the y-intercept of the tangent line to y=(1.2)/(3+7x)^-1/2 at (3,0.245)

OpenStudy (anonymous):

well, i assume you've found the point it is tangent, yes?

OpenStudy (anonymous):

then just create a formula for a line by y = mx + b where m is the slope, and (x,y) is the point it is tangent then solve for b and there ya go!

OpenStudy (anonymous):

Thanks and should I derive thos equation, plug in (3) for X to find m?

OpenStudy (anonymous):

yes, the derivative of the equation at the point (3,.0245) will tell you the slope

OpenStudy (anonymous):

using the given point, the slope you're about to find, you can then solve for b using y = mx + b

OpenStudy (anonymous):

I am having an issue solving for x in this equation.

OpenStudy (anonymous):

you don't need to solve for x

OpenStudy (anonymous):

to find the slope at the tangent point 1. take the derivative 2. plug in known (x,y) values 3. the answer is the slope

OpenStudy (anonymous):

Thanks.....I knew I was missing something simple.

OpenStudy (anonymous):

and you don't have to plug in the y value, by the way just the x value my bad

OpenStudy (anonymous):

you'll be left with something like dy/dx = "number" that number being the slope :)

OpenStudy (anonymous):

i've gotta leave, so to re-cap: 1. find derivative 2. plug in x value 3. you'll be left with something like dy/dx = slope 4. using the given (x,y) point, and the slope you just found 5. plug them into y = mx+b and find b

OpenStudy (anonymous):

good luck mate

OpenStudy (anonymous):

1.derive the equation 2.plug in x 3.the answer is the slope 4.plug the slope, y and x to solve for b 5. let x=0 to find the y-intercept

OpenStudy (anonymous):

Thanks